equations with 3 linear variables
6x-4y+5z=31
5x+2y+2z=13
x+y+z=2
should you multiply the 2nd equation by 2?
show work

Answer by venugopalramana(3286)

OK..THE PROBLEM,WE FACE IN ANSWERING IS THAT WE ARE NOT AWARE OF YOUR COURSE OF STUDY AND BACKGROUND.THE SOLUTION I GAVE IS BY GAUSS JORDAN ELIMINATION METHOD USED IN MATRICES AS WANTED BY SOME STUDENTS.IF YOU ARE NOT TAUGHT THAT ,WE SHALL DO BY NORMAL ELIMINATION.
equations with 3 linear variables
6x-4y+5z=31.......................................I
5x+2y+2z=13...........................II
x+y+z=2.......................III
EQN.II - 2*EQN.III
5X+2Y+2Z-2X-2Y-2Z=13-2*2=9
3X=9
X=9/3=3
EQN.I+4*EQN.III
6X-4Y+5Z+4X+4Y+4Z=31+4*2=39
10X+9Z=39...BUT...X=3..SO
10*3+9Z=39
9Z=39-30=9
Z=9/9=1
PUTTING IN EQN.III
3+Y+1=2
Y=2-4=-2
HOPE YOU UNDERSTOOD.
------------------------------------------------------------------------
SEE THE FOLLOWING AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK
YOU MAY MULTIPLY EQN.II WITH 2 AND ADD TO EQN.I TO ELIMINATE Y
x - y + z = 2
-x + y + z = 4
-x + z = 2
augmented matrix is
1.....................-1.......................1.............................2
-1.....................1.......................1.............................4
-1.....................0.......................1.............................2
nr2=r2+r1 and nr3=r3+r1
1..................-1.................1.............2
1-1=0............1-1=0............1+1=2..........2+4=6
1-1=0...........-1+0=-1...........1+1=2..........2+2=4
nr2=r2/2....nr3=r3-r2
1....................-1...............1..............2
0/2=0.............0/2=0..............2/2=1........6/2=3
0-0=0............-1-0=-1..............2-2=0.......4-6=-2
nr1=r1-r3-r2.......nr2=2r2........nr3=-r3
1-0-0=1............-1+1-0=0.........1-0-1=0.......2+2-3=1
0.....................0...................1............3
0.....................1...................0.............2
exchange r2 and r3
1....0........0.......1
0....1........0.......2
0....0........1.......3
he
nce x=1......y=2.....z=3